Probability and Statistics I GSIA, Carnegie Mellon University 45-733, Spring 2002

Homework #4, due Thursday, February 14, 2002

Please show all relevant work! Please do the following problems from the book:

pp 237-43 3,7,19 pp 250-4 39,45,57

If you get stuck, it would be wise to try to work even-numbered problems similar to the ones assigned, since many of these have solutions in the back of the book.